# Scalable Radar-based ITS: Self-localization and Occupancy Heat Map for Traffic Analysis

### v1

@@ IV-A Localization of Radar Sensors in Real World
For the real world localization, radar data is gathered in Ingolstadt, Germany with a mobile sensor system [1]. It contains a Continental ARS548 long range radar sensor and a Nvidia Orin embedded system to gather and process the data. The radar spans the azimuth range of 120° and the elevation range of 30° and operates at 20Hz. Each frame delivers about 500 detections in a point cloud format. The aerial laser scan is retrieved from open data provided by the Agency for Digitisation, High-Speed Internet and Surveying of Bavaria [28].
@@
A qualitative assessment is presented in Fig. 2. The localization accuracy is demonstrated using the tree canopies as a reference on a parking lot. The radar data was gathered from 23,950 frames over a period of approximately 20 minutes on a windy day. The point clouds first pass through a velocity filter. Only points with a speed higher than $0.15m/s$ are kept. Subsequently, the DBSCAN algorithm is used to filter the cloud with ($\epsilon=0.5,n_{min}=10$). The points at the canopies of trees represent the reflection from the rustling leaves. It is evident that the points are grouped in clusters with a clear separation between each other. The clusters overlap very well with the laser scan point clouds of the trees. This shows that the algorithm accurately determines the position and orientation.
Figure 3: Quantitative evaluation of the localization. The hexagon is the GNSS measured sensor position. 10 points are randomly selected around the ground truth value of the sensor setup as the initial position for the ICP. The error between the ICP result and the ground truth is given in the table.
@@
For a quantitative evaluation, we place the sensor setup at a random position around a junction to gather data in order to perform the localization. Then we compare the results with a ground truth, which is determined with a GNSS device (ArduSimple 2B) using RTK (real-time kinematics) correction. The accuracy of the GNSS measurement is $2cm$. 10 positions in the map surrounding the test spot as initial values for the ICP algorithm. The actual sensor location and the initial starting points are marked in the Fig. 3. We show the 2D errors within the image. The test runs have similar height outcomes, with an average error of 1.81m. Although the lowest 2D position error is $36cm$, the average position error is 1.08m, which is comparable to other map-based approaches [7, 8, 11]. However, we have observed that the point may move far away from its initial position, even when it is located near the actual ground truth. This is likely due to a hidden pattern in the point cloud data. The altitude error may be due to changes in terrain elevation. In Fig. 4 we see that the vehicle’s tracks align with the driving lane while also leaving the green area, road divider, and other areas unoccupied (upper part). It further depicts that the left-turn vehicles stay correctly in the lane (bottom part). This indicates a acceptable accuracy for traffic analysis.
@@
Figure 4: Alignment of the radar point cloud to the road. The lane description is created manually based on an aerial image (dashed lines). The lane information is also used for the masking process in Fig. 1(b). The upper zoomed-in image shows that the green area on the road matches the non-traveled area in the trajectory point cloud. The lower image shows that the traces of the left-turning vehicles match the lane description.
@@ IV-B Localization of Radar Sensors in Carla Simulation
We evaluate the localization method in Carla simulator’s Town 10 map. A laser scan for this map is not directly available. Instead, we convert the road mesh into a point cloud. We then crop the road section using the HD map of the town to remove the parking lanes. We conduct two sets of experiments that change the location and direction of the radar sensors. In the first set of tests, we place the radar sensor with different yaw angles (DY). In the second set of tests, sensors are located at different positions and have a more flexible orientation (DO). In both sets of tests, the sensors are placed as shown in Fig. 5 The sensors have a maximum sensing range of 200m. Like the radar in the real world, the azimuth angle of the sensors covers from -60° to 60°. Their orientations in the first set are also shown in Fig. 5. The elevation range is 30°. The frame rate is 44Hz.
@@
Table I shows the result of the experiments. The 2D location error is less than 1 meter, except for R2 and R6 in DO. The altitude estimate is always off by about 1.06 meters on average. The average yaw errors are 0.38° (DY) and 0.75° (DO). In the most favorable scenario, it could only be off by 0.02° (R1 in DY). The roll and pitch errors in DO are 0.28° and 0.33°, respectively. This outperforms other map-based methods [7, 8, 11]. We believe that the favorable outcome of Carla’s 2D position and orientation is due to its complex observable geometry and simple road geometry. The evaluation demonstrates the potential of our approach.
@@ V Scalable Dynamic Occupancy Map
Figure 6: The global occupancy heat map as a tool for traffic analysis. The heat map is created with data from two sensors. Sensor 1 monitors the road section from top to bottom. Sensor 2 monitors the left-right road section. The maximum range is about 300 meters. The intersection can be monitored by both sensors. The time difference is indicated below the figures. The dark color indicates a high occupancy. The number above the color scale is the maximum occupancy of the polygon nodes in the time scope. For example, 225 implies the occupancy of 11.25s out of 100s ($225\times 50$ms).
After locating multiple radar-based ITS setups in the city (Fig. 1(c)), large-scale traffic analysis could be conducted. It performs better than sensors installed at specific points, such as inductive circuits or ultrasound detectors, within the range. In comparison to vision-based systems, the radar-based systems generate less data traffic while still being able to monitor larger areas using fewer sensors. Using an occupancy heat map as a tool, we can represent the current traffic status and perform spatial statistical analyses.
@@ V-A Using the HD Map as Scaffold
We design the occupancy heat map based on a sublane level HD map. We define the urban occupancy heat map as $M=\{P\}$, where $P$ are the nodes of locations. Unlike scene representation with voxelization, each node in this map is defined as a sub-lane level tile of the road. It consists of a four-vertex polygon with a width of approximately 3m, equal to the street’s width and a length of roughly 0.5m, close to a pedestrian’s standard footstep. These shapes come from a HD map of lanes in Lanelet2 (OSM) format. First, we interpolate the left and right edges of the lanelet to get finer points. Then, every four points that enclose a smallest rectangle form a finer lanelet and are saved as a polygon. A unique ID is assigned to these polygon nodes as a reference. Similar to lanelets at junctions, it is possible to have overlapping nodes.
@@ V-B Local Sensor Data Processing
To analyze the occupancy of the scene, we first map the sensor data to the polygons. Whenever the radar sensor delivers a point cloud, we check their footprints on the east-north plane and assign them to the corresponding polygons. Then, we add the current time stamp to the polygon node’s "occupancy time" attribute, indicating that the node is occupied at that time stamp. Because there may be a very large number of potentially occupied polygons, the R-tree spatial indexing [29] is used for efficient assignment. This compresses the local sensor data into a list of indices of occupied polygons, which can greatly reduce the data volume. Note that the data can be filtered based on the Doppler measurement, the road clearance profile or the radar cross section measurement (reflection strength).
@@ V-C Global Occupancy Heat Map
After connecting sensors to a network, we can generate a global occupancy heat map by fusing all data frames into a single scene representation. We first use NTP (Network-Time-Protocol) to synchronize the time of all data processing units. A 5G module is used to connect the sensor and its data processing unit with the internet. The sensors operate at a nearly constant frequency. However, fusing data from sensors into an integrated map is not as trivial as adding up all the lists of occupied polygons and marking the occupancy with a time stamp. The sensors’ observation areas sometimes overlap and add redundancy. As a result, the overlap areas receive more frequent observations and consequently exhibit higher occupancy. This occurrence suggests incorrectly more occupants in the corresponding areas. To prevent issues, it is necessary to establish a uniform update rate for the global occupancy map that matches the radar update rate. Depending on whether the global occupancy map is generated in a field data processing unit or in the cloud, the update rate can either adopt the frame rate of the connected sensor or be defined independently. Despite efforts to synchronize clocks in data processing units, there is a noticeable time offset in sensor data depending on their starting time (see Fig. 7).
@@
Figure 7: A diagram of message arrival time and time synchronization in the occupancy map. X axis shows the time stamp when messages arrive. Y axis shows the time step between two adjacent messages: $t_{n}=t_{n-1}+y$. The blue rectangles represents sampling windows of 50ms.
@@
The data updates locally at a frequency of approximately 20Hz, but deviations may occur based on the time that onboard algorithms in the sensor may consume. The data for generating the global occupancy heat map is then collected within each 50ms time window. Fig. 7(a) illustrates that if the frame rate is close to 20Hz, we can collect the data and build an occupancy map within each sampling window. If the data frequency deviates significantly (see Fig. 7(b)), two messages may arrive during the same sampling period, resulting in both messages being added to that time step. If no message is received during a particular time step, it should be left empty.
@@
The heat map of occupancy can either be instantly generated to show the current road occupancy status or be cumulatively generated by counting the reported occupancy time stamps to indicate the statistical congestion status. Data volume can be controlled by the number of polygons. Fig. 6 shows an example where two sensors are deployed in the city and localized using the proposed method. At the junction, traffic changes are detected from left to right depending on the occupancy status. Congestion can be observed on the right side of the map, which is due to an increased volume of traffic split into the right turn lane. Similarly, the left-turn waiting area at the intersection from below is congested, indicating a high volume of traffic moving from the upper to the lower portion of the area.
@@ VI Conclusion
In this work, we investigate 4D radar-based ITS to leverage the benefits of long sensing range under diverse weather condition. A method for self-localization is proposed. It leverages idea of matching the concept of "road" in two data modalities: aerial laser scan and live radar sensor data. On one hand, road sections are represented by a point cloud from an aerial laser scan, which is extracted with the help of aerial images and HD maps. On the other hand, the radar point cloud of motion trajectories generated by the observed vehicles also describes the geometry of road sections. These two point clouds are matched with ICP algorithm. The output transformation matrix delivers solutions to the localization task. A extra Kalman Filter can be integrated into the localization process, making corrections during deployments possible. Our results outperforms the result from other map-based radar localization works, especially for the orientation. Using the localization results, we demonstrate the creation of global occupancy heat maps by synchronizing networked radar-based sensor setups for scalable traffic analysis. Based on this work, we contribute to the flexible deployment of radar-based ITS for traffic analysis. The further work lies in detecting individual moving objects on the road and matching their trajectories with the corresponding lane to increase the localization accuracy. Another expansion is adding object information to the occupancy heat map for further traffic analysis.

### v2

@@ IV-A Localization of Radar Sensors in Real World
For real-world localization, radar data is gathered in Ingolstadt, Germany with a mobile sensor system [1]. It contains a Continental ARS548 long-range radar sensor and a Nvidia Orin embedded system to collect and process the data. The radar has an azimuth range of 120° and an elevation range of 30° and operates at 20Hz. Each frame provides about 500 detections in a point cloud format. The aerial laser scan is retrieved from open data provided by the Agency for Digitisation, High-Speed Internet and Surveying of Bavaria [27].
@@
A qualitative assessment is shown in Fig. 2. The localization accuracy is demonstrated using the tree canopies as a reference in a parking lot. The radar data was gathered from 23,950 frames over a period of approximately 20 minutes on a windy day. The point clouds are first passed through a velocity filter. Only points with a speed greater than $0.15m/s$ are retained. The cloud is then filtered using the DBSCAN algorithm (with $\epsilon=0.5,n_{min}=10$). The points on the canopy represent the reflection from the rustling leaves. It can be seen that the points are grouped in clusters with a clear separation between them. The clusters overlap very well with the laser scan point clouds of the trees. This shows that the algorithm is accurate in determining position and orientation.
@@
For a quantitative evaluation, we place the sensor setup at a random position around an intersection to gather data for localization. We then compare the results with a ground truth determined with a GNSS device (ArduSimple 2B) using RTK (real-time kinematics) correction. The accuracy of the GNSS measurement is $2cm$. 50 positions in the map surrounding the test spot are generated as initial values for the ICP algorithm. The actual sensor location, the results, and the initial starting points are marked in the Fig. 3. The 2D errors are also shown. The test runs have similar height results, with an average error of 0.59m. The average 2D position error is 1.06m, which is comparable to other map-based approaches [7, 8, 11], some results of which are also shown in the figure along with the used dataset. The resulted pose are densely distributed around a point, indicating convergence. It also indicates a hidden pattern in the point cloud data. The height error may be due to changes in terrain elevation. The yaw error is extremely low compared to the other methods. This is very important for ITS roadside perception, as we are aiming for long-range detections (up to 300m) with a single sensor, large angular deviations will cause the detection to drift completely away from its true position.
Figure 3: Quantitative evaluation of the localization. 50 initial poses are randomly selected around the ground truth value of the sensor setup for the ICP process. The error between the ICP result and the ground truth is given in the table.
In Fig. 4 we see that the vehicle’s tracks are aligned with the driving lane, while also leaving the green area, road divider, and other areas unoccupied (upper part). It further depicts that the left-turning vehicles stay correctly in the lane (lower part). This indicates an acceptable accuracy for traffic analysis.
@@
Figure 4: Alignment of the radar point cloud to the road. The lane description is manually created based on an aerial image (dashed lines). The lane information is also used for the masking process in Fig. 1(b). The upper zoomed image shows that the green area on the road corresponds to the untraveled area in the trajectory point cloud. The lower image shows that the traces of the left-turning vehicles match the lane description.
@@ IV-B Localization of Radar Sensors in Carla Simulation
We evaluate the localization method in Carla simulator’s Town 10 map. A laser scan is not directly available for this map. Instead, we convert the road mesh into a point cloud. We then crop the road section using the HD map of the town to remove the parking lanes. We run two sets of experiments that change the location and direction of the radar sensors. In the first set of experiments, we place the radar sensor with different yaw angles (DY). In the second set of tests, the sensors are placed at different positions and have a more flexible orientation (DO). In both sets of tests, the sensors are placed as shown in Fig. 5 The sensors have a maximum sensing range of 200m. As with the real-world radar, the azimuth angle of the sensors ranges from -60° to 60°. Their orientations in the first set are also shown in Fig. 5. The elevation range is 30°.
@@
Table I shows the results of the experiments. The 2D position error is less than 1 meter, except for R2 and R6 in DO. The altitude estimate is always off by about 1.06 meters on average. The average yaw errors are 0.38° (DY) and 0.75° (DO). In the best case scenario, it could only be off by 0.02° (R1 in DY). The roll and pitch errors in DO are 0.28° and 0.33° respectively. This also outperforms other map-based methods [7, 8, 11]. We believe that the favorable result of Carla’s 2D position and orientation is due to its complex observable geometry and simple road geometry. The evaluation demonstrates the potential of our approach.
@@ V Scalable Dynamic Occupancy Map
Figure 6: The global occupancy heat map as a tool for traffic analysis. The heat map is created using data from two sensors. Sensor 1 monitors the road section from top to bottom. Sensor 2 monitors the road section from right to left. The maximum range is about 300 meters. The intersection can be monitored by both sensors. The time difference is indicated below the figures. The dark color indicates a high occupancy. The number above the color scale is the maximum occupancy of the polygon nodes in the time scope. For example, 225 implies an occupancy of 11.25s out of 100s ($225\times 50$ms).
After locating multiple radar-based ITS setups in the city (Fig. 1(c)), large-scale traffic analysis could be conducted. It performs better than sensors installed at specific points, such as inductive circuits or ultrasonic detectors, within the area. Compared to vision-based systems, the radar-based systems generate less data traffic and can monitor larger areas with fewer sensors. Using an occupancy heat map as a tool, we can display the current traffic status and perform spatial statistical analysis.
@@ V-A Using the HD Map as Scaffold
We design the occupancy heat map based on a sublane level HD map. We define the urban occupancy heat map as $M=\{P\}$, where $P$ are the nodes of the locations. Unlike the voxelized scene representation, each node in this map is defined as a sub-lane level tile of the road. It is a four-vertex polygon with a width of approximately 3m, equal to the width of the lane and a length of about 0.5m, close to the standard step of a pedestrian. These shapes come from an HD map of lanes in Lanelet2 (OSM) format. First, we interpolate the left and right edges of the lanelet to get finer points. Then, every four points that enclose a smallest rectangle form a finer lanelet and are stored as a polygon. A unique ID is assigned to these polygon nodes as a reference. Similarly to lanelets at junctions, it is possible to have overlapping nodes.
@@ V-B Local Sensor Data Processing
To analyze the occupancy of the scene, we first project the sensor data onto the polygons. Whenever the radar sensor delivers a point cloud, we check its footprints on the east-north plane and assign them to the corresponding polygons. We then add the current timestamp to the polygon node’s "occupancy time" attribute, indicating that the node is occupied at that timestamp. As there can be a very large number of potentially occupied polygons, the R-tree spatial indexing [29] is used for efficient assignment. This compresses the local sensor data into a list of indices of occupied polygons, which can greatly reduce the data volume. Note that the data can be filtered based on the Doppler measurement, the road clearance profile or the radar cross section measurement.
@@ V-C Global Occupancy Heat Map
After connecting the sensors to a network, we can create a global occupancy heat map by merging all the data frames into a single scene representation. We first use NTP (Network Time Protocol) to synchronize the time of all data processing units. A 5G module is used to connect the sensor and its data processing unit to the internet. The sensors operate at a approximately constant frequency. However, fusing sensor data into an integrated map is not as trivial as adding up all the lists of occupied polygons and time stamping the occupancy. Sensor coverage areas sometimes overlap, creating redundancy. As a result, the overlapping areas receive more frequent observations and therefore exhibit a higher occupancy. This will incorrectly suggest more occupants in those areas. To prevent problems, it is necessary to establish a uniform update rate for the global occupancy map that matches the radar update rate. Depending on whether the global occupancy map is generated in a field data processing unit or in the cloud, the update rate can either adopt the frame rate of the connected sensor or be defined independently. Despite efforts to synchronize the clocks in the data processing units, there is a noticeable time offset in the sensor data depending on their starting time (see Fig. 7).
@@
Figure 7: A diagram of message arrival time and time synchronization in the occupancy map. The X-axis shows the time stamp when messages arrive. The Y-axis shows the time step between two adjacent messages: $t_{n}=t_{n-1}+y$. The blue rectangles represent sampling windows of 50ms.
@@
The data is updated locally at a frequency of approximately 20Hz, but there may be variations due to the time taken by the on-board algorithms in the sensor. The data for generating the global occupancy heat map is then collected within each 50ms time window. Fig. 7(a) illustrates that if the frame rate is close to 20Hz, we can collect the data and build an occupancy map within each sampling window. If the data frequency deviates significantly (see Fig. 7(b)), two messages may arrive during the same sampling period, resulting in both messages being added to that time step. If no message is received during a particular time step, it should be left empty.
@@
The Occupancy Heat Map can be generated either instantaneously to show the current state of road occupancy or cumulatively by counting the reported occupancy time stamps to show the statistical state of congestion. The volume of data can be controlled by the number of polygons. Fig. 6 shows an example where two sensors are deployed in the city and localized using the proposed method. The traffic changes at the intersection are shown with the changes in the occupancy status, from left to right. On the right side of the map, congestion can be observed due to an increased volume of traffic in the right turn lane. Similarly, the left-turn queuing area at the intersection from below is congested, indicating a high volume of traffic moving from the bottom to the left part of the area.
@@ VI Conclusion
In this work, we investigate 4D radar-based ITS to leverage the benefits of long sensing range under diverse weather conditions. A method for self-localization is proposed. It exploits the idea of matching the concept of "road" in two data modalities: aerial laser scan and live radar sensor data. On the one hand, road sections are represented by a point cloud from an aerial laser scan, which is extracted with the help of aerial images and HD maps. On the other hand, the radar point cloud of motion trajectories generated by the observed vehicles also describes the geometry of the road sections. These two point clouds are matched using the ICP algorithm. The output transformation matrix delivers solutions to the localization task. An additional Kalman Filter is integrated into the localization process, allowing corrections during deployments. Our results outperform those of other map-based radar localization works, especially for the orientation. Using the localization results, we demonstrate the creation of global occupancy heat maps by synchronizing networked radar-based sensor setups for scalable traffic analysis. Based on this work, we contribute to the flexible deployment of radar-based ITS for traffic analysis. Future work will focus on detecting individual moving objects on the road and matching their trajectories to the corresponding lane to furthur improve the localization accuracy. Another extension is to add object information to the occupancy heat map for further traffic analysis.